# DistJoin: A Decoupled Join Cardinality Estimator based on Adaptive Neural Predicate Modulation

### v2

@@ IV-B DistJoin’s Equi Join Cardinality Inference
Under the assumptions in LABEL:assump:error_distribution and  LABEL:assump:DistJoinAnalysis, this approach provides unbiased estimates with variance proportional to $\text{card}_{J(\mathcal{T})}^{2}$, as proven in subsection IV-C.
@@
Here, denominator vectors follow consistent distinct value encoding and are recursively calculated using the same formula with $\mathcal{Q}=\emptyset$. The algorithm appears in LABEL:alg:InferenceForJoinTypes. DistJoin also supports outer joins by assigning probability 1 to non-existent join key values in corresponding tables before performing equi-joins.
@@ IV-C Expectation and Variance Analysis
We proved the Theorem 2 with the LABEL:assump:DistJoinAnalysis in the Appendix.
@@
    • FactorJoin-style approaches achieve unbiased estimation under standard independence assumptions (LABEL:assump:error_distribution)
    • DistJoin requires stronger regularity conditions specified in LABEL:assump:error_distribution and LABEL:assump:DistJoinAnalysis
@@
    Cardinality Growth Dynamics: Assuming the coefficient term dominated by variance in LABEL:equ:FactorJoinVariance and LABEL:equ:DistJoinVariance is significantly smaller than the baseline value, we have:
@@ IV-D DistJoin’s Non-Equi Join Cardinality Inference
Traversing $v$ from 1 to $|\mathcal{D}(T_{l}.C_{key})|$ enables optimization of repeated sum calculations through the cumsum() function applied to $\hat{\mathbf{P}}_{T_{r}}(T_{r}.C_{key},\mathcal{Q})$. For join conditions like <<, we right-shift the distribution vector, pad with zero at the beginning, and drop the last item before performing cumsum(). The recursive inference process for all supported join conditions appears in LABEL:alg:InferenceForJoinTypes.
@@ V-B Adaptive Neural Predicate Modulation
ANPM’s calculation for $C_{i}^{j}$ follows these formulas, with distinct embedding approaches for training and inference. As shown in LABEL:equ:concat_training, ANPM concatenates factorized true value embedding vectors during training. During inference, ANPM generates combinations of the first j1j-1 sub-columns’ distinct values along the batch dimension, avoiding dimensional explosion as shown in LABEL:equ:concat_inference.
@@ VI-A Experimental Setting
Table III and LABEL:fig:workload_dist reveal that the $10$ test workloads cover extensive ranges of query true cardinality and join schema cardinality. Notably, JOB-light-ranges’ non-equi joins exhibit approximately 4 orders of magnitude lower selectivity, presenting greater prediction challenges. Non-equi joins demonstrate significantly higher true cardinality, with the six tables’ Cartesian product cardinality reaching $1.5e46$ - 13 orders of magnitude above the maximum $\text{card}_{J(\mathcal{T})}$. Consistent with subsection IV-C’s findings, we anticipate DistJoin’s superior performance over FactorJoin. The JOB-light workload with non-equi operators shows substantially higher true cardinality ranges due to simpler predicates, effectively demonstrating query-driven baselines like MSCN’s numerical challenges with non-equi joins. For query-driven baselines, we used ALECE’s workload generation method and generated 100k queries that contain both equi and non-equi join queries.
@@ VI-B Accuracy Evaluation
    DistJoin and ALECE exhibits low bias. As shown in Figure 7, DistJoin and ALECE show significantly lower bias compared to other methods. Although FactorJoin theoretically has a low bias, its binning technique amplifies bias by violating LABEL:assump:error_distribution.

### v3

@@ IV-B DistJoin’s Equi Join Cardinality Inference
Under the assumptions in Assumption 1 and  Assumption 2, this approach provides unbiased estimates with variance proportional to $\text{card}_{J(\mathcal{T})}^{2}$, as proven in subsection IV-C.
@@
Here, denominator vectors follow consistent distinct value encoding and are recursively calculated using the same formula with $\mathcal{Q}=\emptyset$. The algorithm appears in algorithm 2. DistJoin also supports outer joins by assigning probability 1 to non-existent join key values in corresponding tables before performing equi-joins.
@@ IV-C Expectation and Variance Analysis
We proved the Theorem 2 with the Assumption 2 in the Appendix.
@@
    • FactorJoin-style approaches achieve unbiased estimation under standard independence assumptions (Assumption 1)
    • DistJoin requires stronger regularity conditions specified in Assumption 1 and Assumption 2
@@
    Cardinality Growth Dynamics: Assuming the coefficient term dominated by variance in Equation 41 and Equation 63 is significantly smaller than the baseline value, we have:
@@ IV-D DistJoin’s Non-Equi Join Cardinality Inference
Traversing $v$ from 1 to $|\mathcal{D}(T_{l}.C_{key})|$ enables optimization of repeated sum calculations through the cumsum() function applied to $\hat{\mathbf{P}}_{T_{r}}(T_{r}.C_{key},\mathcal{Q})$. For join conditions like <<, we right-shift the distribution vector, pad with zero at the beginning, and drop the last item before performing cumsum(). The recursive inference process for all supported join conditions appears in algorithm 2.
@@ V-B Adaptive Neural Predicate Modulation
ANPM’s calculation for $C_{i}^{j}$ follows these formulas, with distinct embedding approaches for training and inference. As shown in Equation 20, ANPM concatenates factorized true value embedding vectors during training. During inference, ANPM generates combinations of the first j1j-1 sub-columns’ distinct values along the batch dimension, avoiding dimensional explosion as shown in Equation 21.
@@ VI-A Experimental Setting
Table III and Figure 9 reveal that the $10$ test workloads cover extensive ranges of query true cardinality and join schema cardinality. Notably, JOB-light-ranges’ non-equi joins exhibit approximately 4 orders of magnitude lower selectivity, presenting greater prediction challenges. Non-equi joins demonstrate significantly higher true cardinality, with the six tables’ Cartesian product cardinality reaching $1.5e46$ - 13 orders of magnitude above the maximum $\text{card}_{J(\mathcal{T})}$. Consistent with subsection IV-C’s findings, we anticipate DistJoin’s superior performance over FactorJoin. The JOB-light workload with non-equi operators shows substantially higher true cardinality ranges due to simpler predicates, effectively demonstrating query-driven baselines like MSCN’s numerical challenges with non-equi joins. For query-driven baselines, we used ALECE’s workload generation method and generated 100k queries that contain both equi and non-equi join queries.
@@ VI-B Accuracy Evaluation
    DistJoin and ALECE exhibits low bias. As shown in Figure 7, DistJoin and ALECE show significantly lower bias compared to other methods. Although FactorJoin theoretically has a low bias, its binning technique amplifies bias by violating Assumption 1.
@@ Acknowledgments
Appendix A Appendix
Appendix
ANPM Block Formalized Calculation Process
$\overset{(B,F)}{x}$ $=\overset{(B,E)}{E(v_{i}^{1})}\oplus\overset{(B,E)}{E(v_{i}^{2})}\oplus\cdots\oplus\overset{(B,E)}{E(v_{i}^{j-1})}$ (20)
$\begin{split}\overset{(B,F)}{x}&=\Bigl{(}\dots\Bigl{(}\overset{(|\mathcal{D}(C_{i}^{1})|,1,E)}{E(v_{i}^{1}).U(1)}\oplus\overset{(1,|\mathcal{D}(C_{i}^{1})|,E)}{E(v_{i}^{2}).U(0)}\Bigr{)}.U(1)\\ &\quad\oplus\cdots\Bigr{)}\oplus\overset{(1,|\mathcal{D}(1,C_{i}^{1})|,E)}{E(v_{i}^{j-1}).U(0)}\end{split}$ (21)
, where $F=(j-1)\times E$ and $B=\text{batch\_size}$ during training and $B=\prod_{1}^{j-1}|\mathcal{D}(C_{i}^{j})|$ during inference. Then, the ANPM projects the logits by the following formulas:
$\overset{(B,F,1)}{u_{h}}$ $=U(\text{HyperNet}_{h}^{u}(x),2)$ (22)
$\overset{(B,1,E)}{v_{h}}$ $=U(\text{HyperNet}_{h}^{v}(x),1)$ (23)
$\overset{(B,E)}{b_{h}}$ $=\text{HyperNet}_{h}^{b}(x)$ (24)
$\overset{(B,E,1)}{u_{l}}$ $=U(\text{HyperNet}_{l}^{u}(x),2)$ (25)
$\overset{(B,1,F)}{v_{l}}$ $=U(\text{HyperNet}_{l}^{v}(x),1)$ (26)
$\overset{(B,F)}{b_{l}}$ $=\text{HyperNet}_{l}^{b}(x)$ (27)
$\overset{(B,F,E)}{W_{h}}$ $=u_{h}\times v_{h},\quad\overset{(B,E,F)}{W_{l}}=u_{l}\times v_{l}$ (28)
$\overset{(B,F)}{logits_{i}^{\prime j}}$ $=ReLU(\overset{(B,F)}{logits_{i}^{j}}\times W_{h}+b_{h})\times W_{l}+b_{l}$ (29)
, where $U$ represents unsqueezeunsqueeze function.
Proof of the Expectation and Variance of FactorJoin’s Join Cardinality Inference Methods
Before proving Theorem 1, we establish that neural network probability estimation errors follow an arbitrary distribution $\mathcal{G}$. This represents a relaxed constraint compared to typical assumptions of normal distribution in related works [41, 43, 42]. Since our analysis focuses on join cardinality inference accuracy, we also assume unbiased probability estimates for individual tables, implying $\mathbb{E}[\mathcal{G}]=0$.
Assumption 1.
For convenience and clarity, let the bin depth be 1, thus $bin(v)=v$ and $p_{i}(\cdot)$ is the true probability. We assume the error of the predicted probability follows:
$\hat{p}_{T_{i}}(T_{i}.C_{key}$ $=v,\mathcal{Q}(T_{i}))=p_{i}(v,\mathcal{Q})+\epsilon_{i}(v,\mathcal{Q})$ (30)
$\epsilon_{i}(v,\mathcal{Q})$ $\sim\mathcal{G}(0,\sigma_{i}^{2}(v,\mathcal{Q}))$ (31)
$\hat{p}_{T_{i}}(T_{i}.C_{key}$ $=v)=p_{i}(v)+\epsilon_{i}(v)$ (32)
$\epsilon_{i}(v)$ $\sim\mathcal{G}(0,\sigma_{i}^{2}(v))$ (33)
, where $\mathcal{G}(\mu,\sigma^{2})$ represents an arbitrary distribution with expectation $\mu$ and variance $\sigma^{2}$.
With the Assumption 1, we proved the Theorem 1 as follows:
Proof of Theorem 1. Since each table has an independent estimator, the estimated probability of each table is independent. Thus ,  Equation 5 can be proved by:
$\mathbb{E}[\hat{\text{card}}$] $=\mathbb{E}\left[\sum_{v}\prod_{i=1}^{N}\hat{p}_{T_{i}}(T_{i}.C_{key}=v,\mathcal{Q}(T_{i}))|T_{i}|\right]$ (34)
$=\sum_{v}\prod_{i=1}^{N}|T_{i}|\mathbb{E}\left[\hat{p}_{T_{i}}(T_{i}.C_{key}=v,\mathcal{Q}(T_{i}))\right]$ (35)
$=\sum_{v}\prod_{i=1}^{N}|T_{i}|p_{i}(v)$ (36)
Next, we prove the variance is proportional to $\prod_{i=1}^{n}|T_{i}|^{2}$. Based on the designs that both FactorJoin and DistJoin employ independent models to estimate different tables, there are the following facts:
  1. a)
    Since the prediction errors among different key values of the same table are related: $\text{Cov}(\epsilon(v,\mathcal{Q}),\epsilon(v^{\prime},\mathcal{Q}))=C_{i}^{Q}(v,v^{\prime})$
  2. b)
    Since prediction errors among different tables are independent: $\text{Cov}(\epsilon_{i}(v,\mathcal{Q}),\epsilon_{j}(v^{\prime},\mathcal{Q}))=0,\quad(i\neq j)$
Expanding the  Equation 4 by:
$\hat{\text{card}}$ $=\prod_{i=1}^{n}|T_{i}|\sum_{v}\prod_{i=1}^{n}\hat{p}_{i}(v,\mathcal{Q})$ (37)
$=\prod_{i=1}^{n}|T_{i}|\sum_{v}\prod_{i=1}^{n}\left(p_{i}(v,\mathcal{Q})+\epsilon_{i}(v,\mathcal{Q})\right)$ (38)
$=\prod_{i=1}^{n}|T_{i}|\cdot X$ (39)
$X$ $=\sum_{v}\prod_{i=1}^{n}\left(p_{i}(v,\mathcal{Q})+\epsilon_{i}(v,\mathcal{Q})\right)$ (40)
The variance of estimated cardinality is:
[…]
By $X$ is independent of $|T_{i}|$ since $p_{i}$ is normalized probability
$\propto\left(\prod_{i=1}^{n}|T_{i}|\right)^{2}$ (42)
Proof of the Expectation and Variance of DistJoin’s Join Cardinality Inference Methods
Assumption 2.
For analytical convenience, we assume zero covariance between distribution estimates under different inputs:
$\text{Cov}(f(X_{1}),f(X_{2}))=0,\ \text{where}\ f(\cdot)\ \text{is an estimator}$ (43)
Proof of Theorem 2. We first analyze the expectation and variance of the two-table joins situation:
$\hat{\text{card}}$ $=\text{card}_{J(\mathcal{T})}\frac{\sum_{v}\hat{p}_{T_{l}}(v,\mathcal{Q})\hat{p}_{T_{r}}(v,\mathcal{Q})}{\sum_{v}\hat{p}_{T_{l}}(v)\hat{p}_{T_{r}}(v)}$ (44)
By the  Assumption 1:
$\mathbb{E}[\hat{p}_{T_{l}}(v,\mathcal{Q})]$ $=p_{T_{l}}(v,\mathcal{Q}),\quad\mathbb{E}[\hat{p}_{T_{r}}(v,\mathcal{Q})]=p_{T_{r}}(v,\mathcal{Q})$ (45)
$\mathbb{E}[\hat{p}_{T_{l}}(v)]$ $=p_{T_{l}}(v),\quad\mathbb{E}[\hat{p}_{T_{r}}(v)]=p_{T_{r}}(v)$ (46)
Let:
$\hat{S_{\mathcal{Q}}}$ $=\sum_{v}\hat{p}_{T_{l}}(v,\mathcal{Q})\hat{p}_{T_{r}}(v,\mathcal{Q})$ (47)
$\hat{S}$ $=\sum_{v}\hat{p}_{T_{l}}(v)\hat{p}_{T_{r}}(v)$ (48)
$S_{\mathcal{Q}}$ $=\sum_{v}p_{T_{l}}(v,\mathcal{Q})p_{T_{r}}(v,\mathcal{Q})$ (49)
$S$ $=\sum_{v}p_{T_{l}}(v)p_{T_{r}}(v)$ (50)
$\Delta S_{\mathcal{Q}}$ $=\hat{S_{\mathcal{Q}}}-S_{\mathcal{Q}},\quad\Delta S=\hat{S}-S$ (51)
By using independent trained models for different tables and  Assumption 1:
$\mathbb{E}[\hat{S_{\mathcal{Q}}}$] $=\sum_{v}p_{T_{l}}(v,\mathcal{Q})p_{T_{r}}(v,\mathcal{Q})$ (52)
$\mathbb{E}[\hat{S}$] $=\sum_{v}p_{T_{l}}(v)p_{T_{r}}(v)$ (53)
Taylor expansion of $\frac{\hat{S_{\mathcal{Q}}}}{\hat{S}}$ at $\mathbb{E}[\hat{S}$]:
$\frac{\hat{S_{\mathcal{Q}}}}{\hat{S}}$ $\approx\frac{\hat{S_{\mathcal{Q}}}}{\mathbb{E}[\hat{S}]}-\frac{\hat{S_{\mathcal{Q}}}(\hat{S}-\mathbb{E}[\hat{S}])}{\mathbb{E}[\hat{S}]^{2}}+\dots$ (54)
$\mathbb{E}\left[\frac{\hat{S_{\mathcal{Q}}}}{\hat{S}}\right$] $\approx\frac{\mathbb{E}[\hat{S_{\mathcal{Q}}}]}{\mathbb{E}[\hat{S}]}-\frac{\text{Cov}(\hat{S_{\mathcal{Q}}},\hat{S})}{\mathbb{E}[\hat{S}]^{2}}+\dots$ (55)
By  Assumption 2, independent estimation for different tables and ignore the higher-order terms
$\approx\frac{\mathbb{E}[\hat{S_{\mathcal{Q}}}]}{\mathbb{E}[\hat{S}]}$ (56)
$=\frac{\sum_{v}p_{T_{l}}(v,\mathcal{Q})p_{T_{r}}(v,\mathcal{Q})}{\sum_{v}p_{T_{l}}(v)p_{T_{r}}(v)}$ (57)
$\mathbb{E}[\hat{\text{card}}$] $=\text{card}_{J(\mathcal{T})}\mathbb{E}[\frac{\hat{S_{\mathcal{Q}}}}{\hat{S}}$] (58)
$\approx\text{card}_{J(\mathcal{T})}\frac{\sum_{v}p_{T_{l}}(v,\mathcal{Q})p_{T_{r}}(v,\mathcal{Q})}{\sum_{v}p_{T_{l}}(v)p_{T_{r}}(v)}$ (59)
$=\text{card}$ (60)
By $\text{card}_{J(\mathcal{T})}$ is a fixed value that is calculated based on the true distribution, so it’s independent of either $\hat{S_{\mathcal{Q}}}$ or $S$:
[…]
By Taylor expansion and ignoring higher-order terms:
$\frac{\hat{S_{\mathcal{Q}}}}{\hat{S}}$ $\approx\frac{S_{\mathcal{Q}}}{S}+\frac{\Delta S_{\mathcal{Q}}}{S}-\frac{S_{\mathcal{Q}}}{S^{2}}\Delta S$ (62)
$\begin{split}\text{Var}(\hat{\text{card}})&=\text{card}_{J(\mathcal{T})}^{2}\text{Var}\left(\frac{\hat{S_{\mathcal{Q}}}}{\hat{S}}\right)\\ &\approx\text{card}_{J(\mathcal{T})}^{2}\left[\frac{\text{Var}(\Delta S_{\mathcal{Q}})}{S^{2}}+\frac{S_{\mathcal{Q}}^{2}}{S^{4}}\text{Var}(\Delta S)\right.\\ &\quad\left.-2\frac{S_{\mathcal{Q}}}{S^{3}}\text{Cov}(\Delta S_{\mathcal{Q}},\Delta S)\right]\\ &\propto\text{card}_{J(\mathcal{T})}^{2}\end{split}$ (63)
We can use induction to extend the conclusions above to n-table joins. For $k+1$ table joins situation: $T_{1}\Join\dots\Join T_{k+1}$, we regard it as $\mathcal{T}^{(k)}\Join T_{k+1}$, the cardinality can be calculated by:
[…]
$\text{card}_{J(\mathcal{T}^{(k+1)})}$ $=\sum_{v}\left[\text{card}_{J(\mathcal{T}^{(k)})}\cdot p_{\mathcal{T}^{(k)}}(v)\right]\cdot\left[|T_{k+1}|\cdot p_{T_{k+1}}(v)\right]$ (65)
Since $\text{card}_{J(\mathcal{T}^{(k+1)})}$ is a fixed true value, we only need to prove that $\hat{p}_{\mathcal{T}^{(k)}}$ follows  Assumption 1. We proceed by mathematical induction:
Base Case ($k=1$): Based on  Assumption 1:
$\hat{p}_{\mathcal{T}^{(1)}}(v,\mathcal{Q})$ $=p_{\mathcal{T}^{(1)}}(v,\mathcal{Q})+\epsilon_{\mathcal{T}^{(1)}}(v,\mathcal{Q})$ (66)
$\epsilon_{\mathcal{T}^{(1)}}(v,\mathcal{Q})$ $\sim\mathcal{G}^{(1)}(0,\sigma_{1}^{2}(v,\mathcal{Q}))$ (67)
$\hat{p}_{\mathcal{T}^{(1)}}(v)$ $=p_{\mathcal{T}^{(1)}}(v)+\epsilon_{\mathcal{T}^{(1)}}(v)$ (68)
$\epsilon_{\mathcal{T}^{(1)}}(v)$ $\sim\mathcal{G}^{(1)}(0,\sigma_{1}^{2}(v))$ (69)
Inductive Hypothesis: Assuming the k1k-1-th join holds:
$\hat{p}_{\mathcal{T}^{(k-1)}}(v,\mathcal{Q})$ $=p_{\mathcal{T}^{(k-1)}}(v,\mathcal{Q})+\epsilon_{\mathcal{T}^{(k-1)}}(v,\mathcal{Q})$ (70)
$\epsilon_{\mathcal{T}^{(k-1)}}(v,\mathcal{Q})$ $\sim\mathcal{G}^{(k-1)}(0,\sigma_{k-1}^{2}(v,\mathcal{Q}))$ (71)
$\hat{p}_{\mathcal{T}^{(k-1)}}(v)$ $=p_{\mathcal{T}^{(k-1)}}(v)+\epsilon_{\mathcal{T}^{(k-1)}}(v)$ (72)
$\epsilon_{\mathcal{T}^{(k-1)}}(v)$ $\sim\mathcal{G}^{(k-1)}(0,\sigma_{k-1}^{2}(v))$ (73)
Inductive Step: We need to prove that for the $k$-th join, it holds:
$\hat{p}_{\mathcal{T}^{(k)}}(v,\mathcal{Q})$ $=p_{\mathcal{T}^{(k)}}(v,\mathcal{Q})+\epsilon_{\mathcal{T}^{(k)}}(v,\mathcal{Q})$ (74)
$\epsilon_{\mathcal{T}^{(k)}}(v,\mathcal{Q})$ $\sim\mathcal{G}^{(k)}(0,\sigma_{k}^{2}(v,\mathcal{Q}))$ (75)
$\hat{p}_{\mathcal{T}^{(k)}}(v)$ $=p_{\mathcal{T}^{(k)}}(v)+\epsilon_{\mathcal{T}^{(k)}}(v)$ (76)
$\epsilon_{\mathcal{T}^{(k)}}(v)$ $\sim\mathcal{G}^{(k)}(0,\sigma_{k}^{2}(v))$ (77)
Since in  Assumption 1, $\mathcal{G}^{(k)}$ is an arbitrary distribution and $\sigma_{k}^{2}$ is an arbitrary variance, we only need to prove that the expectation of the error $\epsilon_{\mathcal{T}^{(k)}}$ is zero.
By  Equation 10, we have:
$\hat{p}_{\mathcal{T}^{(k)}}(v,\mathcal{Q})$ $=\frac{\hat{p}_{\mathcal{T}^{(k-1)}}\left(v,\mathcal{Q}\left(\mathcal{T}^{(k-1)}\right)\right)\cdot\hat{p}_{T_{k}}\left(v,\mathcal{Q}\left(T_{k}\right)\right)}{\sum_{v}\hat{p}_{\mathcal{T}^{(k-1)}}(v)\hat{p}_{T_{k}}(v)}$ (78)
Let:
$\hat{S}_{Q}^{(k)}$ $=\hat{p}_{\mathcal{T}^{(k-1)}}\left(v,\mathcal{Q}\left(\mathcal{T}^{(k-1)}\right)\right)\cdot\hat{p}_{T_{k}}\left(v,\mathcal{Q}\left(T_{k}\right)\right)$ (79)
$\hat{S}^{(k)}$ $=\sum_{v}\hat{p}_{\mathcal{T}^{(k-1)}}(v)\hat{p}_{T_{k}}(v)$ (80)
$S_{\mathcal{Q}}^{(k)}$ $=p_{\mathcal{T}^{(k-1)}}\left(v,\mathcal{Q}\left(\mathcal{T}^{(k-1)}\right)\right)\cdot p_{T_{k}}\left(v,\mathcal{Q}\left(T_{k}\right)\right)$ (81)
$S^{(k)}$ $=\sum_{v}p_{\mathcal{T}^{(k-1)}}(v)p_{T_{k}}(v)$ (82)
$\Delta S_{\mathcal{Q}}^{(k)}$ $=\hat{S}_{Q}^{(k)}-S_{\mathcal{Q}}^{(k)},\qquad\Delta S^{(k)}=\hat{S}^{(k)}-S^{(k)}$ (83)
Similar to the formula Equation 55, by Taylor expansion, we can be obtained:
$\mathbb{E}\left[\hat{p}_{\mathcal{T}^{(k)}}(v,\mathcal{Q})\right]$ $\approx\frac{\mathbb{E}\left[\hat{S_{\mathcal{Q}}}^{(k)}\right]}{\mathbb{E}\left[\hat{S}^{(k)}\right]}-\frac{\text{Cov}\left(\hat{S_{\mathcal{Q}}}^{(k)},\hat{S}^{(k)}\right)}{\mathbb{E}\left[\hat{S}^{(k)}\right]^{2}}+\dots$ (84)
By  Assumption 2 and independent estimation between different tables:
$\approx\frac{\mathbb{E}\left[\hat{S_{\mathcal{Q}}}^{(k)}\right]}{\mathbb{E}\left[\hat{S}^{(k)}\right]}$ (85)
By using independent trained models for different tables and the inductive hypothesis:
$=\frac{S_{\mathcal{Q}}^{(k)}}{S^{(k)}}$ (86)
Because :
$\mathbb{E}\left[\hat{p}_{\mathcal{T}^{(k)}}(v,\mathcal{Q})\right]$ $=\mathbb{E}\left[p_{\mathcal{T}^{(k)}}(v,\mathcal{Q})+\epsilon_{\mathcal{T}^{(k)}}(v,\mathcal{Q})\right]$ (87)
$=p_{\mathcal{T}^{(k)}}(v,\mathcal{Q})+\mathbb{E}\left[\epsilon_{\mathcal{T}^{(k)}}(v,\mathcal{Q})\right]$ (88)
$=\frac{S_{\mathcal{Q}}^{(k)}}{S^{(k)}}+\mathbb{E}\left[\epsilon_{\mathcal{T}^{(k)}}(v,\mathcal{Q})\right]$ (89)
We can obtain that:
$\mathbb{E}\left[\epsilon_{\mathcal{T}^{(k)}}(v,\mathcal{Q})\right]$ $=0$ (90)
End of Induction Proof. Thus, because the base case holds and the inductive step holds, we proved that for $N$-table joins, the $\epsilon_{\mathcal{T}^{(k)}}(v,\mathcal{Q})\sim\mathcal{G}^{(k)}(0,\sigma_{k}^{2}(v,\mathcal{Q}))$ holds. By the same reasoning, let $\mathcal{Q}=\emptyset$, the $\epsilon_{\mathcal{T}^{(k)}}(v)\sim\mathcal{G}^{(k)}(0,\sigma_{k}^{2}(v))$ holds. That is, $\hat{p}_{\mathcal{T}^{(k)}}\left(v,\mathcal{Q}\left(\mathcal{T}^{(k)}\right)\right)$ follows the  Assumption 1.
Now we use induction to extend the conclusions of two-table joins to N-table joins:
Base Case ($k=2$):
$\mathbb{E}[\hat{\text{card}}^{(2)}]$ $\approx\text{card}_{J(\mathcal{T}^{(2)})}\frac{\sum_{v}p_{T_{l}}(v,\mathcal{Q})p_{T_{r}}(v,\mathcal{Q})}{\sum_{v}p_{T_{l}}(v)p_{T_{r}}(v)}$ (91)
$=\text{card}^{(2)}$ (92)
$\text{Var}(\hat{\text{card}}^{(2)})$ $\propto\text{card}_{J(\mathcal{T}^{(2)})}^{2}$ (93)
Inductive Hypothesis: Assuming the $k$-table joins hold:
$\mathbb{E}[\hat{\text{card}}^{(k)}]$ $\approx\text{card}_{J(\mathcal{T}^{(k)})}\frac{\sum_{v}p_{\mathcal{T}^{(k-1)}}(v,\mathcal{Q})p_{T_{k}}(v,\mathcal{Q})}{\sum_{v}p_{\mathcal{T}^{(k-1)}}(v)p_{T_{k}}(v)}$ (94)
$=\text{card}^{(k)}$ (95)
$\text{Var}(\hat{\text{card}}^{(k)})$ $\propto\text{card}_{J(\mathcal{T}^{(k)})}^{2}$ (96)
Inductive Step: Based on  Equation 64,  Assumption 2, and the conclusion we just proved that $\hat{p}_{\mathcal{T}^{(k)}}\left(v,\mathcal{Q}\left(\mathcal{T}^{(k)}\right)\right)$ follows the  Assumption 1, we can regard $\mathcal{T}^{(k-1)}$ as an individual table. Then, with the same reason that we proved the base case, the following formulas hold for $n=k$ situation.
$\mathbb{E}[\hat{\text{card}}^{(k)}]$ $\approx\text{card}_{J(\mathcal{T}^{(k)})}\frac{\sum_{v}p_{\mathcal{T}^{(k-1)}}(v,\mathcal{Q})p_{T_{k}}(v,\mathcal{Q})}{\sum_{v}p_{\mathcal{T}^{(k-1)}}(v)p_{T_{k}}(v)}$ (97)
$=\text{card}^{(k)}$ (98)
$\text{Var}(\hat{\text{card}}^{(k)})$ $\propto\text{card}_{J(\mathcal{T}^{(k)})}^{2}$ (99)
End of Proof. In summary, we first proved that  Theorem 2 holds for the two-table joins situation, and then we used induction and proved that the conclusion can be extended to the n-table joins situation. Thus, we have proven  Theorem 2. ∎
Algorithm of how DistJoin deals with non-equi joins
Input: probability distributions: $\hat{\mathbf{P}}$
Output: $\hat{\mathbf{P}}_{\mathcal{T}}(C_{key},\mathcal{Q})$
1 $p_{l}(\mathcal{Q}),p_{r}(\mathcal{Q}),p_{l},p_{r}\longleftarrow\hat{\mathbf{P}}_{\mathcal{T}_{left}}(C_{key},\mathcal{Q}),\hat{\mathbf{P}}_{\mathcal{T}_{right}}(C_{key},\mathcal{Q}),\hat{\mathbf{P}}_{\mathcal{T}_{left}}(C_{key}),\hat{\mathbf{P}}_{\mathcal{T}_{right}}(C_{key})$;
2 if $join\_how=^{\prime}\geq^{\prime}$ then
3  $p_{l}(\mathcal{Q})\longleftarrow p_{l}(\mathcal{Q})+\sum p_{l}(\mathcal{Q})-\text{cumsum}(p_{l}(\mathcal{Q}))$ ; //reverse cumsum
4
5 if $join\_how=^{\prime}\leq^{\prime}$ then
6  $p_{l}(\mathcal{Q})\longleftarrow\text{cumsum}(p_{l}(\mathcal{Q}))$ ;
7
8 if $join\_how=^{\prime}>^{\prime}$ then
9  $p_{l}(\mathcal{Q})\longleftarrow\text{Roll}(p_{l}(\mathcal{Q}),-1)$ ; //Circularly shift left one position
10  $p_{l}(\mathcal{Q})[-1]\longleftarrow 0;$
11  $p_{l}(\mathcal{Q})\longleftarrow p_{l}(\mathcal{Q})+\sum p_{l}(\mathcal{Q})-\text{cumsum}(p_{l}(\mathcal{Q}))$ ; //reverse cumsum
12
13 if $join\_how=^{\prime}<^{\prime}$ then
14  $p_{l}(\mathcal{Q})\longleftarrow\text{Roll}(p_{l}(\mathcal{Q}),1)$ ; //Circularly shift right one position
15  $p_{l}(\mathcal{Q})[0]\longleftarrow 0;$
16  $p_{l}(\mathcal{Q})\longleftarrow\text{cumsum}(p_{l}(\mathcal{Q}))$ ;
17
18 $\hat{\mathbf{P}}_{\mathcal{T}}(C_{key},\mathcal{Q})\longleftarrow\frac{p_{l}(\mathcal{Q})\odot p_{r}(\mathcal{Q})}{\sum_{v}p_{l}\cdot p_{r}}$ ;
19
20return $\hat{\mathbf{P}}_{\mathcal{T}}(C_{key},\mathcal{Q})$;
Algorithm 2 Inference of Join Cardinality for Different Join Conditions
Additional Experiment
Supplementary Settings
Model Architectures. Each table’s ANPM model includes:
  • ResMADE with 4 layers (256 units each)
  • ANPM HyperNetworks with 2 layers
  • Factorized value bit width: 12
  • Predicate embedding size: 64
Since the ANPM Blocks are the performance bottleneck, we also list the parameter numbers of each model’s ANPM Blocks in Table VII. The results demonstrate that despite the large #DV values for each table, ANPM’s low-rank architecture ensures the corresponding model maintains a remarkably small number of ANPM parameters.
TABLE VII: Number of ANPM Blocks’ parameters.
Model of table Parameter Numbers
cast_info 49920
movie_companies 99840
movie_info 49920
movie_info_idx 49920
movie_keyword 99840
title 127140
Workload Distribution
The workloads’ distributions are shown as Figure 9
Figure 9: Distribution of each workload. The join conditions $(>,\geq)$ and $(<,\leq)$ have very similar selectivity distribution, so their curves overlay each other.
Comprehensive Q-Error comparison
We demonstrate the comprehensive Q-Error comparison of all join types and more detailed percentiles in Figure 10. The results show that DistJoin outperforms all baselines on almost all Q-Error percentiles.
Comparison on TPC-H datasets
We generated complex join queries with range predicates on the TPC-H dataset to compare the most representative baseline’s accuracy with DistJoin’s. Following Naru’s workload generation method [1], we randomly selected join tables and then generated predicates on each of them. Finally, we calculate the true cardinality of them based on the equi join condition. The results in  Table VIII show that DistJoin also significantly outperforms those baselines.
TABLE VIII: Q-Error comparison on TPC-H
Methods mean 50th 95th 99th max
PG 2.47 1.33 6.22 17.01 82.67
FactorJoin $6.7e11$ 8.57 211.26 $4.0e8$ $6.6e14$
DistJoin 2.19 1.50 5.67 12.50 38.0
Evaluation of Convergence Speed
We illustrate DistJoin’s convergence curve in Figure 11. Since all tables train in parallel, grouping models by epoch proves difficult. Instead, we use the longest training time among all tables per epoch as the x-axis to demonstrate DistJoin’s performance convergence over 20 epochs. DistJoin converges quickly in terms of median metrics, suggesting that for scenarios less sensitive to long-tail error distributions, training time can be reduced by accepting slightly increased long-tail error distributions for faster training.
Figure 11: DistJoin’s convergence curve.